A gas sample occupies a volume of 12.0 L at 25°C and 750 torr. At what temperature in °C will the gas occupy a volume of 15.0 L at 750 torr?
1
Expert's answer
2020-10-16T08:24:52-0400

 volume V1 =12.0 L 

temperature T1 = 25°C = 273 K

 and Pressure = 750 torr.


new volume V2 =15.0 L at 750 torr

temperature T2 = ?


combined gas law says


V1/T1 = V2/T2


T2 = V2 * T1/V1 = 15 L * 273 K/12 L = 341.25 K


Answer


Final temperature = 341 K